How many grams of O2 can you make with 92.46 grams of H2O


1
Expert's answer
2021-03-29T06:05:21-0400

2H2O -> 2H2 +O2

2 moles of H2O = 1 m of O2

Molar Mass of H2O = 18.01528

Molar Mass of O2 = 31.998

= 92.46/18.01528 = 5.1323 moles of H20

= (5.1323/2) × 31.998

= 82.1116677g of O2